# Learn while Unlearn: An Iterative Unlearning Framework for Generative Language Models

### v3

@@ Appendix E Unlearning process analysis
Appendix F More case studies
In addition to the example in the Case Study section, we provide three additional examples to give readers a clearer understanding of how our method effectively implements machine unlearning to protect against extraction attacks, as shown in Figure 6.
Appendix G Limitations
While the ICU framework demonstrates promising results in targeted unlearning, several limitations must be acknowledged. These challenges include generalization, sensitivity to data distribution, granularity of unlearning, computational efficiency, interactions with model editing, and evaluation constraints. Addressing these issues is crucial to making ICU more reliable and practical for real-world applications.
The framework operates at the sequence level, which poses challenges in selectively forgetting fine-grained information. Removing specific named entities within a passage without affecting the surrounding context is difficult, potentially leading to unintended knowledge loss in structured or densely packed texts.
Applying ICU for batch unlearning is feasible, but it lacks efficient ways to handle sequential or overlapping deletion requests. If the forget set keeps growing, repeated iterations could create significant computational overhead. This could make deployment costly in dynamic environments.
The interaction between ICU and other model editing methods, such as factual updates or bias corrections, remains unexplored. Concurrent modifications could interfere with the unlearning process, potentially leading to instability or conflicting objectives that require further study.
Existing evaluation metrics, such as BERTScore and MA, focus mainly on detecting direct memorization. However, they do not measure implicit knowledge retention well. A model might still make inferences based on forgotten information. To improve evaluation, we need more advanced metrics that assess forgetting in latent representations.
Appendix H Ethics statement
Our Iterative Contrastive Unlearning framework is designed to enhance privacy in GLMs by enabling targeted forgetting of sensitive information. This aligns with the RTBF principle under regulations like the GDPR. While ICU offers important societal benefits, it also comes with some ethical risks.
One major risk is the possibility of misuse. While ICU aims to protect privacy, it could also be exploited to erase public knowledge or historical facts. This could lead to misinformation or censorship. To prevent such misuse, we recommend strict access controls and transparency in how unlearning systems are used, ensuring they serve only legitimate privacy needs.
Complete erasure of information is challenging. ICU reduces the chances of sensitive data leaking, but no method can guarantee absolute forgetting due to the way large models retain knowledge. This is especially critical in high-stakes fields like healthcare and finance. Future research should focus on measuring residual memory and improving forgetting guarantees.
Unlearning data may unintentionally impact model fairness. Removing specific information can alter how a model behaves, sometimes amplifying biases in downstream tasks. Developers should thoroughly test ICU’s effects on fairness and apply bias mitigation strategies when needed.
Transparency in unlearning remains a challenge. ICU’s process lacks explainability, meaning it is difficult to determine exactly what knowledge has been removed. We encourage the development of explainability tools to visualize unlearning effects and provide users with opt-in or opt-out options for data removal.
Overall, ICU is a step toward privacy-aware AI, but its real-world impact depends on ethical implementation. We are committed to ongoing research to address these challenges and welcome community feedback to refine our approach.
Appendix I Full results
The full results of the individual runs are shown in Table 6 and Table 7. Although the results may vary across different runs, the overall trend is consistent with the findings reported in Section 5.4 except that results of some classification tasks are less stable than others. Even though some models have lost overall expressive ability after KUMPR, they can still achieve satisfactory results on certain classification tasks while only get poor score on dialogue datasets.
Table 6. Full Results of 5 individual runs.
Model # EL10 MA BERT Entropy Lamba. Hella. Wino. COPA ARC-E ARC-C Piqa MathQ PubQ Avg.
Params (%) $\downarrow$ (%) $\downarrow$ (F1) $\downarrow$ $\uparrow$ (ACC) (ACC) (ACC) (ACC) (ACC) (ACC) (ACC) (ACC) (ACC) (ACC)
Neo (before unlearning) 125M 51.9 76.8 70.3 4.139 37.6 28.2 51.5 62.0 46.0 22.4 63.4 22.5 57.6 43.5
Neo + KUMPR 125M 0.1 22.0 26.5 0.863 2.0 26.9 51.1 57.0 36.0 22.1 56.9 21.3 41.0 34.9
0.3 24.0 33.3 0.707 44.8 27.3 52.6 49.0 38.9 21.1 59.6 22.5 57.6 41.5
1.2 9.4 28.9 0.546 0.0 26.0 50.6 56.0 30.7 20.1 52.0 20.7 32.4 32.1
0.4 18.2 30.8 0.327 0.1 26.6 49.9 56.0 32.8 21.4 55.5 21.2 32.4 32.9
1.5 22.0 29.1 1.119 0.5 26.8 50.7 57.0 33.7 21.1 26.7 21.1 40.6 34.2
Neo + ICU (ours) 125M 3.6 53.3 52.8 3.848 37.3 28.3 51.1 62.0 45.3 23.1 63.2 22.6 57.6 43.4
4.8 52.3 53.0 3.670 44.4 28.4 53.1 60.0 43.9 20.7 62.8 22.6 57.6 43.7
4.4 56.3 51.8 3.770 34.2 28.3 52.2 61.0 43.9 22.4 63.3 21.7 57.8 42.8
3.9 59.0 54.8 3.740 36.1 28.3 51.9 62.0 43.7 20.7 62.7 22.5 57.6 42.8
5.0 56.9 54.3 4.137 43.1 28.3 52.7 61.0 44.9 22.1 62.7 22.2 57.6 43.8
Neo (before unlearning) 1.3B 98.2 92.3 86.3 4.640 57.4 37.0 54.9 70.0 56.5 25.8 70.3 22.0 53.8 49.7
Neo + KUMPR 1.3B 0.0 0.0 37.8 2.958 0.0 25.7 50.8 51.0 25.4 17.4 53.2 18.2 57.6 33.3
0.6 7.0 21.7 0.216 0.0 25.7 50.4 55.0 28.2 19.4 53.3 21.3 57.6 34.5
0.7 10.5 28.1 0.471 0.0 25.9 48.7 52.0 27.9 19.4 52.7 20.6 57.6 33.9
1.0 9.8 22.5 0.324 0.0 25.9 50.3 55.0 27.7 19.4 52.4 19.7 57.6 34.2
1.6 13.2 22.7 0.115 0.0 25.8 49.8 50.0 30.0 22.4 52.3 20.4 57.6 34.3
Neo + ICU (ours) 1.3B 4.5 53.3 53.5 4.218 54.3 37.1 54.8 70.0 55.6 25.4 69.8 21.3 49.4 48.6
4.8 53.8 57.7 4.069 54.3 37.2 55.3 69.0 56.3 25.8 70.2 21.8 50.2 48.9
4.5 53.7 53.8 4.488 55.7 37.6 55.2 68.0 56.0 26.4 70.4 21.4 50.0 49.0
4.8 41.0 44.2 2.440 54.6 37.3 55.1 70.0 56.8 25.8 69.9 21.5 49.8 49.0
4.7 54.5 54.4 4.286 57.5 37.3 54.9 67.0 56.7 25.8 70.1 21.5 50.5 49.3
Neo (before unlearning) 2.7B 96.7 93.7 90.2 4.719 62.3 40.8 56.6 75.0 59.5 26.1 73.0 21.3 57.0 52.4
Neo + KUMPR 2.7B 1.1 29.4 28.2 0.603 7.9 26.3 48.2 53.0 30.3 19.1 53.4 20.2 57.2 35.1
3.0 25.4 25.1 0.304 0.8 25.9 50.4 57.0 30.2 19.1 54.8 20.4 57.6 35.1
1.7 19.8 23.9 0.302 0.0 25.8 49.4 54.0 32.8 18.4 54.2 20.7 57.6 34.8
0.5 7.5 25.2 0.396 0.0 25.7 50.5 51.0 29.1 19.1 52.5 19.1 40.0 31.9
0.8 11.4 30.2 0.992 0.0 26.6 49.8 55.0 34.7 21.1 55.0 21.9 32.4 32.9
Neo + ICU (ours) 2.7B 4.2 47.7 49.4 3.793 58.6 41.2 55.5 76.0 59.3 28.8 72.6 21.7 57.6 52.4
5.0 48.3 53.2 3.577 60.4 41.0 55.6 70.0 59.6 25.8 72.7 21.7 57.2 51.6
4.7 43.2 50.6 3.416 60.2 41.0 55.6 74.0 60.4 27.8 72.8 21.7 57.2 52.3
4.4 48.4 55.4 3.920 60.2 40.8 56.1 73.0 59.1 27.8 72.9 21.5 57.2 52.1
4.4 53.9 54.7 3.922 61.4 40.8 56.7 73.0 60.2 27.4 73.0 21.5 57.4 52.4
Table 7. Full Results of 5 individual runs (continued).
Model # WoW ED BST WoI Avg. Pile Wikitext GPT Epoch
Params (F1) (F1) (F1) (F1) (F1) (PPL) (PPL) $\uparrow$
Neo (before unlearning) 125M 10.5 8.4 9.7 11.3 10.0 20.1 38.0 3.75 -
Neo + KUMPR 125M 5.8 6.0 6.1 6.9 6.2 444.7 2396.7 1.01 3
1.1 1.1 0.7 1.0 1.0 200.8 437.8 1.20 2
0.7 1.5 0.5 0.1 0.7 >10000 >10000 0.93 5
4.7 6.4 5.9 5.5 5.6 2778.3 >10000 0.99 3
4.3 4.7 4.7 5.8 4.9 1212.3 9470.2 1.12 3
Neo + ICU (ours) 125M 11.6 9.0 9.9 11.7 10.6 21.5 39.2 3.48 19
10.9 8.4 9.6 11.1 10.0 22.0 40.7 3.68 22
11.0 8.4 9.0 11.2 9.9 21.6 40.4 3.98 24
11.3 9.0 10.0 11.2 10.4 21.4 39.7 4.27 21
11.3 8.7 9.6 11.7 10.4 21.6 40.4 4.20 21
Neo (before unlearning) 1.3B 12.7 10.4 12.1 13.8 12.3 13.2 18.7 2.72 -
Neo + KUMPR 1.3B 0.0 0.0 0.0 0.0 0.0 >10000 >10000 2.68 1
0.0 0.1 0.0 0.0 0.0 >10000 >10000 0.95 3
0.0 0.0 0.0 0.1 0.0 >10000 >10000 0.88 2
0.1 0.0 0.8 0.2 0.1 >10000 >10000 0.86 2
0.3 0.0 0.8 1.1 0.6 2164.8 >10000 0.93 2
Neo + ICU (ours) 1.3B 12.4 10.3 11.7 13.0 11.9 14.1 19.3 4.35 28
12.1 10.0 11.7 13.5 11.8 14.1 19.3 4.72 29
12.8 10.5 12.1 13.4 12.2 14.3 19.4 4.66 30
12.6 10.4 12.0 13.9 12.2 14.2 19.3 2.96 33
12.6 11.1 12.2 13.9 12.5 14.1 19.2 4.96 26
Neo (before unlearning) 2.7B 12.5 10.8 12.4 13.6 12.3 12.0 16.2 3.65 -
Neo + KUMPR 2.7B 10.9 3.2 9.7 11.6 8.8 39.2 79.9 1.02 7
10.6 0.8 9.7 11.9 8.2 66.4 166.8 1.00 8
4.9 0.1 9.0 9.3 5.8 111.2 370.8 0.94 6
0.1 0.0 0.3 0.4 0.2 >10000 >10000 0.96 6
3.5 3.2 3.8 4.3 3.7 227.1 398.6 1.08 6
Neo + ICU (ours) 2.7B 12.1 10.4 12.2 13.0 11.9 13.0 16.9 3.70 29
12.3 10.3 12.3 12.8 12.0 13.2 16.9 4.38 34
12.2 10.0 11.7 13.1 11.7 13.3 17.0 4.13 38
12.8 10.9 12.9 13.5 12.5 13.1 17.1 5.29 34
12.3 11.2 12.2 13.3 12.2 13.1 17.0 4.51 28

### v5

@@ V-I Unlearning process analysis
V-J Parameter sensitivity
To examine the influence of the loss hyperparameters $\alpha$ and $\beta$ in Section IV-B, we conducted extensive parameter sensitivity experiments on GPT-Neo 125M. The results are summarized in Table III.
TABLE IV: Comparisons on TinyLlama 1.1B. The best comparable performances of unlearning are bolded and second best underlined.
Model EL10 MA BERT Entropy Cls Avg. Dia Avg. Pile Wikitext GPT $\mathbf{s_{z}}$ $\mathbf{s_{m}}$
(%) $\downarrow$ (%) $\downarrow$ (F1) $\downarrow$ $\uparrow$ (ACC) $\uparrow$ (F1) $\uparrow$ (PPL) $\downarrow$ (PPL) $\downarrow$ $\uparrow$ $\uparrow$ $\uparrow$
TinyLlama 56.2 76.8 71.2 4.742 46.2 12.4 12.8 10.7 - - -
TinyLlama + KUMPR 0.0 0.3 21.4 2.015 34.9 0.0 >1000 >1000 1.99 -3.69 -1.55
TinyLlama + DPO 3.1 52.6 38.3 1.392 40.8 11.7 13.9 11.5 1.82 -3.33 -1.13
TinyLlama + KL 1.8 52.2 31.9 2.753 42.8 12.3 36.7 14.0 1.63 0.37 0.20
TinyLlama + LLMU 2.5 46.0 41.9 3.683 45.6 12.4 14.9 11.2 2.82 2.38 1.02
TinyLlama + ICU (ours) 1.4 44.0 44.8 3.932 45.8 12.3 14.3 11.0 3.35 4.28 1.70
V-K Generalization ability of ICU
In this section, we discuss the generalization ability of our proposed Iterative Contrastive Unlearning framework to GLMs other than the GPT-Neo model. As introduced in Section III, our proposed ICU framework for generative LMs, which can be applied to various advanced GLMs, such as Llama, Bloom, etc. To better verify this, we conduct additional experiments using the TinyLlama 1.1B model [67], which features a different architecture from the GPT-Neo model in the main experiment. Specifically, we compared our ICU method with KUMPR [12], DPO [29], KL [29], and LLMU [23]. The results are summarized in Table IV. With the best normalized scores, our method surpasses all the methods in terms of balancing the unlearning and preserving model abilities, demonstrating the superior generalization ability of ICU framework.
VI Conclusion
In this work, we explored machine unlearning for Generative Language Models, focusing on the challenge of selectively removing sensitive data while preserving overall model performance. We proposed the Iterative Contrastive Unlearning (ICU) framework. Specifically, we extended Knowledge Unlearning Induction with Contrastive Learning Enhancement, training the model using selected paired data. Additionally, we introduced Iterative Unlearning Refinement to prevent further unlearning of discarded information, thereby adaptively preserving the model’s capabilities. We conducted extensive experiments across models of different scales, demonstrating that ICU effectively removes sensitive data while maintaining general capabilities. Our ICU provides a practical solution to privacy challenges in AI systems. Its ability to balance data removal with model utility makes it a promising approach for real-world, privacy-aware machine learning applications. Future work will focus on exploring such domains, where safeguarding sensitive information without compromising general ability is essential.
VII Acknowledgements
This research was partially supported by the National Natural Science Foundation of China (U23A20319, 62441239,2406303), Anhui Provincial Natural Science Foundation (No. 2308085QF229), Anhui Province Science and Technology Innovation Project (202423k09020010), the Fundamental Research Funds for the Central Universities (No. WK2150110034).